OpenStudy (anonymous):

For each of the following equilibria, write the equilibrium constant expression for K sub c. (also)? Where appropriate, write the equilibrium constant expression for K sub p. (a) Fe2O3(s) + 3CO(g) <----> 2Fe(l) + 3CO2(g) (b) 4Fe(s) + 3O2(g) <----> 2Fe2O3(s) (c) BaSO4(s) <----> BaO(s) + SO3(g) (d) BaSO4(s) <----> Ba2+(aq) + SO4 2- (aq)

OpenStudy (abb0t):

To write the equlibtrium expression for a reaction in the form: \(\sf \color{red}{_nA+_nB \rightarrow_nC+_nD}\)\(\sf \color{}{K_{p/c} = \frac{[products]}{[reactants]} = \frac{[C]^n[D]^{n}}{[A]^{n}[B]^{n}}}\) Note: "n" represents mole number.